OpenStudy (anonymous):

find xintercepts and vertex of the porabola y=x^2+10x+25

OpenStudy (anonymous):

y=(x+5)^2 ===> vertex = (-5,0) , x-itercept is x=-5

OpenStudy (anonymous):

-5 is root and (-5,0) is the vertex. http://www.wolframalpha.com/input/?i=+y%3Dx^2%2B10x%2B25+vertex

OpenStudy (anonymous):

find the xintercepts and vertext of the porabola y=x^2+2x-24

OpenStudy (ybarrap):

y = (x + 1)^2+ -25 no x-intercept since vertex is at (-1,-25) http://www.wolframalpha.com/input/?i=y+%3D++%28x+%2B+1%29^2%2B+-25+vertex